If you throw exactly two heads in two tosses of a coin you win $75. If not, you pay me $21

step 1.Find the expected value of the proposition. Round your answer to two decimal places. Losses must be expressed as negative values

step 2. If you played this game 700 times how much would you expect to win or lose? Round your answer to two decimal places. Losses must be entered as negative
Answer #1

1)P(two heads in a row)=(1/2)*(1/2)=1/4

expected value of the proposition =75*(1/4)-21*(3/4)=3

2)

expected to win/lose in 700 games =3*700=2100